### v2

# Pólya’s conjecture for thin products

Abstract.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded Euclidean domain. According to the famous Weyl law, both its Dirichlet eigenvalue $\lambda_{k}(\Omega)$ and its Neumann eigenvalue $\mu_{k}(\Omega)$ have the same leading asymptotics $w_{k}(\Omega)=C(d,\Omega)k^{2/d}$ as $k\to\infty$. G. Pólya conjectured in 1954 that each Dirichlet eigenvalue $\lambda_{k}(\Omega)$ is greater than $w_{k}(\Omega)$, while each Neumann eigenvalue $\mu_{k}(\Omega)$ is no more than $w_{k}(\Omega)$. In this paper we prove Pólya’s conjecture for thin products, i.e. domains of the form $(a\Omega_{1})\times\Omega_{2}$, where $\Omega_{1},\Omega_{2}$ are Euclidean domains, and $a$ is small enough. We also prove that the same inequalities hold if $\Omega_{2}$ is replaced by a Riemannian manifold, and thus get Pólya’s conjecture for a class of “thin” Riemannian manifolds with boundary.
Partially supported by National Key R and D Program of China 2020YFA0713100, and by NSFC no. 12171446.
1. Introduction
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain. Then the Dirichlet Laplacian on $\Omega$ has discrete spectrum which forms an increasing sequence of positive numbers (each with finite multiplicity) that tend to infinity,
$0<\lambda_{1}(\Omega)\leq\lambda_{2}(\Omega)\leq\lambda_{3}(\Omega)\leq\cdots\nearrow+\infty,$
and the Neumann Laplacian on $\Omega$ has a similar discrete spectrum (under suitable boundary regularity assumptions, which we always assume below without further mentioning)
$0=\mu_{0}(\Omega)\leq\mu_{1}(\Omega)\leq\mu_{2}(\Omega)\leq\cdots\nearrow+\infty.$
Moreover, by a simple variational argument one has $\mu_{k-1}(\Omega)<\lambda_{k}(\Omega)$ for all $k$, which was strengthened to
(1.1) $\mu_{k}(\Omega)<\lambda_{k}(\Omega),\quad\forall k$
by L. Friedlander in [15] (See also N. Filonov [6]), answering a conjecture of L. E. Payne [27]. Starting from H. Weyl ([35]), the asymptotic behavior of the eigenvalues $\lambda_{k}(\Omega)$ and $\mu_{k}(\Omega)$ as $k\to\infty$ has attracted a lot of attention. In fact, both $\lambda_{k}(\Omega)$ and $\mu_{k}(\Omega)$ admit the same leading term asymptotics
$\lambda_{k}(\Omega)\sim\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{2}{d}}\quad\text{and}\quad\mu_{k}(\Omega)\sim\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{2}{d}},$
where $|\Omega|$ represents the volume of $\Omega$, and $\omega_{d}$ is the volume of the unit ball in $\mathbb{R}^{d}$. In his classical book [29], G. Pólya conjectured (in a slightly weaker form for the Neumann case) that for each $k$, the $k^{\mathrm{th}}$ Dirichlet eigenvalue
(1.2) $\lambda_{k}(\Omega)\geq\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{2}{d}}$
while the $k^{\mathrm{th}}$ positive Neumann eigenvalue
(1.3) $\mu_{k}(\Omega)\leq\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{2}{d}}.$
As observed by G. Pólya, these conjectured inequalities hold for all rectangles. As for arbitrary domain, the conjecture holds for $k=1$ (the Faber-Krahn inequality ([5], [19]) for the Dirichlet eigenvalue, and the Szegö-Weinberger inequality ([33], [34]) for the Neumann case) and $k=2$ (the Krahn-Szegö inequality ([20]) for the Dirichlet case, and recently proved by D. Bucur and A. Henrot in [4] for the Neumann case). The first major progress on the conjecture was made by G. Pólya himself in 1961 ([30]), in which he presented an elegant proof of his conjecture for planar tiling domains (in fact G. Pólya’s proof for the Neumann eigenvalue case was a bit complicated and assumed the tiling to be regular. The regularity assumption was removed and the proof was simplified in 1966 by R. Kellner [17]). Very recently, N. Filonov, M. Levitin, I. Polterovich and D. Sher ([7]) proved that Pólya’s conjecture holds for planar disks (and for Euclidean balls of all dimensions for the Dirichlet case), and thus gave the first non-tiling planar domain for which Pólya’s conjecture is known to be true. For an arbitrary Euclidean domain $\Omega\subset\mathbb{R}^{d}$, P. Li and S.T. Yau proved in [23] that
(1.4) $\sum_{j=1}^{k}\lambda_{k}(\Omega)\geq\frac{d}{d+2}\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{d+2}{d}},$
and as a consequence, got a weaker version of Pólya’s inequality for all Dirichlet eigenvalues,
(1.5) $\lambda_{k}(\Omega)\geq\frac{d}{d+2}\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{2}{d}}.$
P. Kröger proved in [18] a similar weaker upper bound for the Neumann eigenvalues of any Euclidean domain with piecewise smooth boundary. Another important class of domains satisfying Pólya’s conjecture was obtained by A. Laptev [21], in which he proved that if Pólya’s conjecture (1.2) holds for $\Omega_{1}\subset\mathbb{R}^{d_{1}}$, where $d_{1}\geq 2$, then Pólya’s conjecture (1.2) also holds for any domain of the form $\Omega=\Omega_{1}\times\Omega_{2}$. One key ingredient in his proof is the following inequality (which is a special case of Berezin-Lieb inequality ([3], [24]) and is equivalent to Li-Yau’s inequality (1.4) above) for the Riesz mean,
(1.6) $\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}\leq L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}},$
where $\gamma\geq 1$, and
(1.7) $L_{\gamma,d}=\frac{\Gamma(\gamma+1)}{(4\pi)^{\frac{d}{2}}\Gamma(\gamma+1+\frac{d}{2})}.$
For Neumann eigenvalues, A. Laptev also got a similar inequality
(1.8) $\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}\geq L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}$
using which one can get Pólya’s conjecture (1.3) for $\Omega=\Omega_{1}\times\Omega_{2}$ provided $\Omega_{1}$ satisfies (1.3) and has dimension $d_{1}\geq 2$. For other recent progresses concerning Pólya’s conjecture, we refer to [8],[11], [12], [13], [25] etc. In this paper we will prove Pólya’s conjecture for domains of product type that are “thin” in one component, namely regions of the form
$\Omega=a\Omega_{1}\times\Omega_{2}$
for $a$ small enough, without assuming that $\Omega_{1}$ or $\Omega_{2}$ satisfies Pólya’s conjecture. We first prove
Theorem 1.1.
Let $\Omega_{1}\subset\mathbb{R}^{d_{1}}$ and $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ be bounded Euclidean domains, where $d_{1},d_{2}\geq 2$, and $\Omega_{2}$ has piecewise smooth boundary.
  1. (1)
    If $\Omega_{1}$ has Lipschitz boundary, then there exists $a_{0}>0$ (depends on $\Omega_{1}$ and $\Omega_{2}$) such that for any $0<a<a_{0}$, the product $\Omega=a\Omega_{1}\times\Omega_{2}$ satisfies the Dirichlet Pólya’s conjecture (1.2).
  2. (2)
    If $\Omega_{1}$ has $C^{1}$ boundary, then there exists $a_{0}>0$ (depends on $\Omega_{1}$ and $\Omega_{2}$) such that for any $0<a<a_{0}$, the product $\Omega=a\Omega_{1}\times\Omega_{2}$ satisfies the Neumann Pólya’s conjecture (1.3).
Here is the strategy of proof: Following Laptev’s argument [21], we write the eigenvalue counting function of $a\Omega_{1}\times\Omega_{2}$ as the sum of many eigenvalue counting functions of $\Omega_{2}$. Although we don’t have Pólya’s inequality for $\Omega_{2}$, we do have weaker inequalities (See (2.6) and (2.7) below) that follow from Seeley’s version of the two-term Weyl law (which only requires $\Omega_{2}$ to have piecewise smooth boundary). Now instead of applying Laptev’s inequalities on Riesz mean above, we apply stronger two-term inequalities on Riesz mean, namely (2.8) obtained by R. Frank and S. Larson in [10] (for $\Omega_{1}$ to have Lipschitz boundary) and its Neumann analogue, (2.9), obtained by R. Frank and L. Geisinger in [9] (for $\Omega_{1}$ to have $C^{1}$ boundary) to control the sum of both terms in Seeley’s inequalities. By comparing what we lose from Seeley’s two-term bound and what we gain from these two-term Riesz mean bound, we are able to prove that for $a$ small enough, Pólya’s inequalities hold for $\lambda$ large enough (which depends on $a$). For smaller $\lambda$, we use a simple observation that Laptev’s inequalities above are in fact strict (See §2.2 below), and thus (by taking $a$ even smaller) give us the demanded gap to prove Pólya’s inequality. This argument works perfectly well for $d_{2}\geq 3$, but the first part fails for $d_{2}=2$ since in this case we can’t apply Laptev-type inequality on Riesz mean (which requires $\gamma=\frac{d_{2}-1}{2}\geq 1$) to control the sum of the second term of Seeley’s inequality. Fortunately, we can overcome this problem by using Li-Yau’s estimate (1.5) above and an explicit integral computation. The Neumann case is a bit simpler, since we only need Weyl’s law to control the second term in Seeley’s inequality for large $\lambda$, and thus we don’t need to distinguish the case $d_{2}=2$ with $d_{2}\geq 3$.
Note that Laptev’s argument does not work for the case $d_{1}=1$, since the inequalities (1.6) and (1.8) require $\gamma\geq 1$. Even though the interval $(0,1)$ tiles $\mathbb{R}$, it is still not known whether $(0,1)\times\Omega$ satisfies Pólya’s conjecture for general $\Omega$. In the second part of this paper, we turn to study Pólya’s conjecture for thin products $(0,a)\times\Omega$. Instead of writing the eigenvalue counting function of $(0,a)\times\Omega$ as the sum of many eigenvalue counting functions of $(0,a)$ (which is a tiling domain) that we have a nice control, we will write it as the sum of many eigenvalue counting functions of $\Omega$ and apply Seeley’s two-term inequalities. By carefully analyzing the two sums, we shall prove that in this case, all thin products satisfy Pólya’s conjecture:
Theorem 1.2.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded domain with piecewise smooth boundary, then there exists $a_{0}>0$ (depends on $\Omega$) such that for any $0<a<a_{0}$, $(0,a)\times\Omega$ satisfies Pólya’s conjecture (1.2) and (1.3).
Since scaling will not affect Pólya’s inequalities, we immediately see that for any bounded Euclidean domain $\Omega$, there exists a constant $C>0$ such that for all $A>C$, $(0,1)\times A\Omega$ satisfies Pólya’s conjecture. Unfortunately we still can’t prove Pólya’s conjecture for products of the form $(0,1)\times a\Omega$ for small $a$, which obviously implies Pólya’s conjecture for $(0,1)\times\Omega$. The last part of this paper devoted to Pólya’s inequalities for Riemannian manifolds with boundary. Although the original conjecture was only proposed for Euclidean domains, people do study the analogous problem in the more general Riemannian setting. For example, P. Bérard and G. Besson proved in [2] that for a 2-dimensional hemisphere (or a quarter of a sphere, or even an octant of a sphere), both Dirichlet eigenvalues and Neumann eigenvalues satisfy Pólya’s inequalities above. Recently in [14], P. Freitas, J. Mao and I. Salavessa studied the problem for hemispheres in arbitrary dimension. They showed that (1.3) holds for Neumann eigenvalues of hemispheres in any dimension, while (1.2) fails for Dirichlet eigenvalues when $d>2$, and they derived sharp inequality for Dirichlet eigenvalues by adding a correction term. It is thus a natural problem to find out more Riemannian manifolds with boundary satisfying Pólya’s inequalities. Note that in the proof of Theorem 1.1 and Theorem 1.2, for $\Omega_{2}$ and $\Omega$ we mainly used Seeley’s two-term Weyl’s inequality. As a result, by literally repeating the proof one can easily see that for any closed Riemannian manifold $M$, the Neumann eigenvalues of the product $a\Omega\times M$ satisfy Pólya conjecture (1.3) as long as $a$ is small enough. For the Dirichlet case, there will be one extra term (since 00 is an eigenvalue of $M$) in the eigenvalue counting function of the product, namely the number of eigenvalues of $\Omega$ that is less than $a^{2}\lambda$, which can be explicitly calculated if $d_{1}=\dim\Omega=1$ and can be controlled via Li-Yau’s estimate (1.5) if $d_{2}\geq 2$. As a result, we are able to prove that Pólya’s conjecture holds for such Riemannian manifolds with boundary:
Theorem 1.3.
Let $\Omega\subset\mathbb{R}^{d_{1}}$ be a bounded domain with $C^{1}$ boundary and $(M,g)$ be a closed Riemannian manifold of dimension $d_{2}\geq 2$. Then there exists $a_{0}>0$ (depends on $\Omega$ and $M$) such that for any $0<a<a_{0}$, $a\Omega\times M$ satisfies Pólya’s conjecture (1.2) and (1.3).
The arrangement of this paper is as follows. In Section 2 we will list the two-term inequalities for the eigenvalues counting functions and for the Riesz means that will be used later, and also prove strict version of Laptev’s inequalities for completeness. In Section 3 we will prove Theorem 1.1, and in Section 4 we will prove Theorem 1.2. In Section 5 we will turn to the Riemannian manifold setting and prove Theorem 1.3. Moreover we will explain how to get similar results for a larger class of eigenvalue problems. Finally in Section 6 we will give an explicit non-tiling planar domain $\Omega$ and explicitly calculate the constant involved in the proof, and as a result, show that the Dirichlet eigenvalues of $[0,\frac{1}{4\pi}]\times\Omega$ for that $\Omega$ satisfies (1.2). Similarly in the Riemannian setting we show that $(0,a)\times S^{2}$ satisfies Pólya’s conjecture for $a\leq\frac{\pi}{24}$, but will break both Pólya’s inequalities if $a$ is large.
2. Some preparations
2.1. Two term inequalities for the eigenvalue counting functions and the Riesz means
For any bounded domain $\Omega\subset\mathbb{R}^{d}$, we denote the Dirichlet eigenvalue counting function by
$\mathcal{N}^{D}_{\Omega}(\lambda):=\#\{n:\ \lambda_{n}(\Omega)<\lambda\},$
and the Neumann eigenvalue counting function by
$\mathcal{N}^{N}_{\Omega}(\lambda):=\#\{n:\ \mu_{n}(\Omega)<\lambda\}.$
Then the inequality (1.1) implies
$\mathcal{N}^{D}_{\Omega}(\lambda)\leq\mathcal{N}^{N}_{\Omega}(\lambda),\quad\forall\lambda>0,$
while Pólya’s conjectures (1.2) and (1.3) can be restated as
(2.1) $\mathcal{N}^{D}_{\Omega}(\lambda)\leq C_{d}|\Omega|\lambda^{\frac{d}{2}},\qquad\forall\lambda>0,$
for all bounded domains, and
(2.2) $\mathcal{N}^{N}_{\Omega}(\lambda)\geq C_{d}|\Omega|\lambda^{\frac{d}{2}},\qquad\forall\lambda>0,$
for all bounded domains with suitable boundary regularity, where the constant
(2.3) $C_{d}=\frac{\omega_{d}}{(2\pi)^{d}}=\frac{1}{(4\pi)^{\frac{d}{2}}\Gamma(\frac{d}{2}+1)}=L_{0,d}.$
Since the unit balls satisfy $B^{d_{1}+d_{2}}\subset B^{d_{1}}\times B^{d_{2}}$, one has $\omega_{d_{1}+d_{2}}<\omega_{d_{1}}\cdot\omega_{d_{2}}$ and thus
(2.4) $C_{d_{1}+d_{2}}<C_{d_{1}}\cdot C_{d_{2}}.$
It was first obtained by H. Weyl ([35]) that both eigenvalue counting functions $\mathcal{N}^{D}_{\Omega}(\lambda)$ and $\mathcal{N}^{N}_{\Omega}(\lambda)$ have the same leading asymptotics
(2.5) $\mathcal{N}^{D/N}_{\Omega}(\lambda)=C_{d}|\Omega|\lambda^{\frac{d}{2}}+o(\lambda^{\frac{d}{2}})$
as $\lambda\to\infty$, and the famous Weyl’s conjecture, proven by V. Ivrii ([16]) and R. Melrose ([26]) under extra assumptions on the behavior of billiard dynamics, claims that for $\Omega\subset\mathbb{R}^{d}$ with piecewise smooth boundary,
$\mathcal{N}^{D}_{\Omega}(\lambda)=C_{d}|\Omega|\lambda^{\frac{d}{2}}-\frac{1}{4}C_{d-1}|\partial\Omega|\lambda^{\frac{d-1}{2}}+o(\lambda^{\frac{d-1}{2}})$
while
$\mathcal{N}^{N}_{\Omega}(\lambda)=C_{d}|\Omega|\lambda^{\frac{d}{2}}+\frac{1}{4}C_{d-1}|\partial\Omega|\lambda^{\frac{d-1}{2}}+o(\lambda^{\frac{d-1}{2}}),$
where $|\partial\Omega|$ is the surface area of $\partial\Omega$. Although Weyl’s conjecture was not proven in its full generality, R. Seeley ([31], [32]) proved a weaker version, namely both eigenvalue counting functions satisfy
$\mathcal{N}^{D/N}_{\Omega}(\lambda)=C_{d}|\Omega|\lambda^{\frac{d}{2}}+\mathrm{O}(\lambda^{\frac{d-1}{2}}),\qquad\text{as\ }\lambda\to\infty,$
for all bounded domains in $\mathbb{R}^{d}$ with piecewise smooth boundary. In view of the facts $\lambda_{1}(\Omega)>0$ and $\mu_{0}(\Omega)=0$, we see that there exists a positive constant $C(\Omega)$ such that for any $\lambda>0$,
(2.6) $\mathcal{N}^{D}_{\Omega}(\lambda)\leq C_{d}|\Omega|\lambda^{\frac{d}{2}}+C(\Omega)\lambda^{\frac{d-1}{2}}$
and
(2.7) $\mathcal{N}^{N}_{\Omega}(\lambda)\geq C_{d}|\Omega|\lambda^{\frac{d}{2}}-C(\Omega)\lambda^{\frac{d-1}{2}}.$
These two-term inequalities sharpen Weyl’s leading estimates and will play a crucial role below.
We also need two-term inequalities for the Riesz mean that sharpen Laptev’s inequalities (1.6) and (1.8). For the Dirichlet case, R. L. Frank and S. Larson ([10]) proved that for any bounded domain $\Omega$ in $\mathbb{R}^{d}$ ($d\geq 2$) with Lipschitz boundary and any $\gamma\geq 1$,
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}=L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\frac{1}{4}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}+\mathrm{o}(\lambda^{\gamma+\frac{d-1}{2}})$
as $\lambda\to\infty$. As a consequence, for fixed $\gamma$, there exists a positive constant $C_{1}(\Omega)$ such that if $\lambda>C_{1}(\Omega)$, one has
(2.8) $\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}\leq L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\frac{1}{5}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}.$
For the Neumann case, R. L. Frank and L. Geisinger ([9]) proved that for any bounded domain $\Omega$ in $\mathbb{R}^{d}$ ($d\geq 2$) with $C^{1}$ boundary and any $\gamma\geq 1$, one has
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}=L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}+\frac{1}{4}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}+\mathrm{o}(\lambda^{\gamma+\frac{d-1}{2}})$
as $\lambda\to\infty$. As a consequence, for fixed $\gamma$, there exists a positive constant $C_{2}(\Omega)$ such that if $\lambda>C_{2}(\Omega)$, one has
(2.9) $\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}\geq L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}+\frac{1}{5}L_{\gamma,d-1}|\partial\Omega|\lambda^{\gamma+\frac{d-1}{2}}.$
2.2. Strict inequalities for Riesz means
Another ingredient in our proof is the inequality (1.6) of A. Laptev, which follows from
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))\leq L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
By carefully analyzing his proof, one can see that the inequality is strict. We include the proof for reader’s convenience.
Proposition 2.1.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{d}$, then for any $\lambda>0$, one has
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))<L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
Proof.
We first sketch Laptev’s proof ([21, Theorem 2.1]). Let $\{\varphi_{k}\}_{k=1}^{\infty}$ be the $L^{2}$ normalized eigenfunctions associated to $\{\lambda_{k}(\Omega)\}_{k=1}^{\infty}$. Then
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))\leq(2\pi)^{-d}\int_{\mathbb{R}^{d}}(\lambda-|\xi|^{2})_{+}\cdot\sum_{\lambda_{k}(\Omega)<\lambda}|\hat{\varphi}_{k}(\xi)|^{2}\mathrm{d}\xi.$
Let
$e_{\xi}(x)=\begin{cases}e^{-ix\cdot\xi},&\qquad x\in\Omega,\\ 0,&\qquad x\notin\Omega.\end{cases}$
Then
$\sum_{\lambda_{k}(\Omega)<\lambda}|\hat{\varphi}_{k}(\xi)|^{2}=\sum_{\lambda_{k}(\Omega)<\lambda}\langle e_{\xi},\varphi_{k}\rangle^{2}_{L^{2}(\Omega)}\\ \leq\|e_{\xi}\|^{2}_{L^{2}(\Omega)}=|\Omega|.$
So
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))\leq(2\pi)^{-d}|\Omega|\int_{\mathbb{R}^{d}}(\lambda-|\xi|^{2})_{+}\mathrm{d}\xi=L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
Now assume by contradiction that $\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))=L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}$ for some $\lambda$, then
$\sum_{\lambda_{k}(\Omega)<\lambda}\langle e_{\xi},\varphi_{k}\rangle^{2}_{L^{2}(\Omega)}=\|e_{\xi}\|^{2}_{L^{2}(\Omega)},\qquad\forall|\xi|^{2}<\lambda.$
Take $N$ such that $\lambda_{N}(\Omega)<\lambda\leq\lambda_{N+1}(\Omega)$, then one has $e_{\xi}\in\mathrm{span}\{\varphi_{1},\cdots,\varphi_{N}\}$ for any $\xi$ with $|\xi|^{2}<\lambda$. So if we take $N+1$ different points $\{\xi_{1},\cdots,\xi_{N+1}\}$ with $|\xi_{i}|^{2}<\lambda$, $\forall 1\leq i\leq N+1$, then there exists $(c_{1},\cdots,c_{N+1})\neq(0,\cdots,0)$ such that
$\sum_{i=1}^{N+1}c_{i}e_{\xi_{i}}(x)=0,\qquad\forall x\in\Omega.$
By the analyticity of $e^{-ix\cdot\xi}$, one has
$\sum_{i=1}^{N+1}c_{i}e^{-ix\cdot\xi_{i}}=0,\qquad\forall x\in\mathbb{R}^{d},$
which is a contradiction. ∎
Together with the fact
(2.10) $\int^{\infty}_{0}(z-\lambda-t)_{+}t^{\gamma-2}\mathrm{d}t=\frac{(z-\lambda)^{\gamma}_{+}}{\gamma(\gamma-1)},\qquad\forall\gamma>1,$
we immediately get
Corollary 2.2.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{d}$, then for any $\lambda>0$ and $\gamma\geq 1$, one has
$\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}<L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}.$
In particular, for any $0<a<b$, one has
(2.11) $K(a,b,\Omega):=\inf_{a\leq\lambda\leq b}\frac{L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}-\sum_{\lambda_{k}(\Omega)<\lambda}(\lambda-\lambda_{k}(\Omega))^{\gamma}}{\lambda^{\gamma+\frac{d-1}{2}}}>0.$
We also need the Neumann version of the same (strict) inequality, which is essentially proved by A. Laptev (c.f. [21], Theorem 3.1):
Proposition 2.3.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{d}$ with discrete Neumann spectrum. Then for any $\lambda\geq 0$, one has
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))>L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
Proof.
Again we first sketch Laptev’s proof of the corresponding (non-strict) inequality in [21]. Let $\{\psi_{k}\}_{k=0}^{\infty}$ be the $L^{2}$ normalized eigenfunctions associated to $\{\mu_{k}(\Omega)\}_{k=0}^{\infty}$ and denote $\varphi_{\lambda}(x)=(\lambda-x)_{+}$. Then
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))=(2\pi)^{-d}|\Omega|\int_{\mathbb{R}^{d}}\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))\mathrm{d}\xi.$
Using the facts
$\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}=1\quad\text{and}\quad\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}=|\xi|^{2}$
and the Jensen inequality (since $\varphi_{\lambda}$ is convex), Laptev proved
$\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))\geq\varphi_{\lambda}(|\xi|^{2})$
and thus
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))\geq L_{1,d}|\Omega|\lambda^{\frac{d+2}{2}}.$
To show that the inequality is strict, again we fix $\lambda$ and let $N$ be the integer with $\mu_{N}(\Omega)\leq\lambda<\mu_{N+1}(\Omega)$. Take $\xi$ with $|\xi|^{2}<\lambda$ such that
$\langle e_{\xi},1\rangle_{L^{2}(\Omega)}\neq 0,\text{ and }e_{\xi}\notin\mathrm{span}\{\psi_{1},\cdots,\psi_{N}\},$
then for such a $\xi$, Jensen’s inequality is strict:
$\sum_{k=0}^{\infty}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))$
$=$ $\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))+\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\varphi_{\lambda}(\mu_{k}(\Omega))$
$=$ $(\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)})\cdot\varphi_{\lambda}(\frac{\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}(\Omega)}{\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}})+$
$(\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)})\cdot\varphi_{\lambda}(\frac{\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}(\Omega)}{\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}})$
$>$ $\varphi_{\lambda}(\sum_{\mu_{k}(\Omega)<\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}(\Omega)+\sum_{\mu_{k}(\Omega)>\lambda}\frac{1}{|\Omega|}\langle e_{\xi},\psi_{k}\rangle^{2}_{L^{2}(\Omega)}\cdot\mu_{k}(\Omega))$
$=$ $\varphi_{\lambda}(|\xi|^{2}).$
This completes the proof. ∎
As a consequence,
Corollary 2.4.
Let $\Omega$ be a bounded domain in $\mathbb{R}^{d}$ with discrete Neumann spectrum. Then for any $\lambda\geq 0$ and $\gamma\geq 1$, one has
$\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}>L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}.$
In particular, for any $0<a<b$, one has
$K_{1}(a,b,\Omega):=\inf_{a\leq\lambda\leq b}\frac{\sum_{\mu_{k}(\Omega)<\lambda}(\lambda-\mu_{k}(\Omega))^{\gamma}-L_{\gamma,d}|\Omega|\lambda^{\gamma+\frac{d}{2}}}{\lambda^{\gamma+\frac{d-1}{2}}}>0.$
3. Proof of Theorem 1.1
As observed by P. Freitas, J. Lagace and J. Payette in [12, Proposition 3.1], it is enough to assume that both $\Omega_{1}$ and $\Omega_{2}$ are connected. We divide the proof of Theorem 1.1 into three parts: the Dirichlet case with $d_{2}\geq 3$, the Dirichlet case with $d_{2}=2$, and the Neumann case. For the Dirichlet case, the eigenvalues of $a\Omega_{1}\times\Omega_{2}$ are
$a^{-2}\lambda_{l}(\Omega_{1})+\lambda_{k}(\Omega_{2}),\qquad\forall l,k\in\mathbb{Z}_{>0}$
and thus
$\mathcal{N}^{D}_{a\Omega_{1}\times\Omega_{2}}(\lambda)=\sum_{l=1}^{Z^{\lambda}_{a}}\mathcal{N}^{D}_{\Omega_{2}}(\lambda-a^{-2}\lambda_{l}(\Omega_{1})),$
where
$Z^{\lambda}_{a}=\mathcal{N}^{D}_{\Omega_{1}}(a^{2}\lambda).$
By inequality (2.6), there exists a constant $C(\Omega_{2})>0$ such that
$\mathcal{N}^{D}_{\Omega_{2}}(\lambda)\leq C_{d_{2}}|\Omega_{2}|\lambda^{\frac{d_{2}}{2}}+C(\Omega_{2})\lambda^{\frac{d_{2}-1}{2}},\qquad\forall\lambda>0.$
So we get
(3.1) $\sum_{l=1}^{Z^{\lambda}_{a}}\mathcal{N}^{D}_{\Omega_{2}}(\lambda-a^{-2}\lambda_{l}(\Omega_{1}))$
$\leq$ $C_{d_{2}}|\Omega_{2}|\sum_{l=1}^{Z^{\lambda}_{a}}(\lambda-a^{-2}\lambda_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}+C(\Omega_{2})\sum_{l=1}^{Z^{\lambda}_{a}}(\lambda-a^{-2}\lambda_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}$
$=$ $C_{d_{2}}|\Omega_{2}|a^{-d_{2}}\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}+C(\Omega_{2})a^{1-d_{2}}\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}.$
By inequality (2.8), there exists a constant $C(\Omega_{1})>0$ such that if $a^{2}\lambda>C(\Omega_{1})$, then
(3.2) $\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}\leq L_{\frac{d_{2}}{2},d_{1}}|\Omega_{1}|a^{d_{1}+d_{2}}\lambda^{\frac{d_{1}+d_{2}}{2}}-\frac{1}{5}L_{\frac{d_{2}}{2},d_{1}-1}|\partial\Omega_{1}|a^{d_{1}+d_{2}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
3.1. The Dirichlet case with $d_{2}\geq 3$
By Corollary 2.2, one has
(3.3) $\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}<L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|a^{d_{1}+d_{2}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
So by (3.1), (3.2), (3.3) and the fact
$C_{d_{2}}L_{\frac{d_{2}}{2},d_{1}}=C_{d_{1}+d_{2}},$
one has that if $a^{2}\lambda>C(\Omega_{1})$, then
$\sum_{l=1}^{Z^{\lambda}_{a}}\mathcal{N}^{D}_{\Omega_{2}}(\lambda-a^{-2}\lambda_{l}(\Omega_{1}))$
$\leq$ $C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}-\frac{1}{5}C_{d_{1}+d_{2}-1}|\partial\Omega_{1}||\Omega_{2}|a^{d_{1}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}$
$+L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})a^{d_{1}}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
Thus if we assume
$a<\frac{C_{d_{1}+d_{2}-1}|\partial\Omega_{1}||\Omega_{2}|}{5L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})}=\frac{C_{d_{2}-1}|\partial\Omega_{1}||\Omega_{2}|}{5|\Omega_{1}|C(\Omega_{2})},$
then for any $\lambda>a^{-2}C(\Omega_{1})$, we will get the demanded inequality
$\mathcal{N}^{D}_{a\Omega_{1}\times\Omega_{2}}(\lambda)\leq C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}.$
Note that if $0<\lambda<a^{-2}\lambda_{1}(\Omega_{1})$, then we automatically have
$\mathcal{N}^{D}_{a\Omega_{1}\times\Omega_{2}}(\lambda)=0<C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}.$
So it remains to consider the case $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C(\Omega_{1})$ assuming $C(\Omega_{1})>\lambda_{1}(\Omega_{1})$. Let $\mu=a^{2}\lambda$, then by Corollary 2.2, one has
$K(\Omega_{1})=\inf_{\lambda_{1}(\Omega_{1})\leq\mu\leq C(\Omega_{1})}\frac{L_{\frac{d_{2}}{2},d_{1}}|\Omega_{1}|\mu^{\frac{d_{1}+d_{2}}{2}}-\sum_{\lambda_{l}(\Omega_{1})<\mu}(\mu-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}}{\mu^{\frac{d_{1}+d_{2}-1}{2}}}>0$
and thus
(3.4) $\sum_{\lambda_{l}(\Omega_{1})<\mu}(\mu-\lambda_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}\leq L_{\frac{d_{2}}{2},d_{1}}|\Omega_{1}|\mu^{\frac{d_{1}+d_{2}}{2}}-K(\Omega_{1})\mu^{\frac{d_{1}+d_{2}-1}{2}}$
for all $\lambda_{1}(\Omega_{1})\leq\mu\leq C(\Omega_{1})$. Thus by (3.1), (3.3) and (3.4), one has that if $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C(\Omega_{1})$, then
$\sum_{l=1}^{Z^{\lambda}_{a}}\mathcal{N}^{D}_{\Omega_{2}}(\lambda-a^{-2}\lambda_{l}(\Omega_{1}))$
$\leq$ $C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}-K(\Omega_{1})C_{d_{2}}|\Omega_{2}|a^{d_{1}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}$
$+L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})a^{d_{1}}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
So if we assume $a<\frac{K(\Omega_{1})C_{d_{2}}|\Omega_{2}|}{L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})}$, then for any $a^{-2}\lambda_{1}(\Omega_{1})\leq\lambda\leq a^{-2}C(\Omega_{1})$,
$\mathcal{N}^{D}_{a\Omega_{1}\times\Omega_{2}}(\lambda)\leq C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}.$
Combining all discussions above, one has that if
$a<\min\bigg{(}\frac{C_{d_{2}-1}|\partial\Omega_{1}||\Omega_{2}|}{5|\Omega_{1}|C(\Omega_{2})},\frac{K(\Omega_{1})C_{d_{2}}|\Omega_{2}|}{L_{\frac{d_{2}-1}{2},d_{1}}|\Omega_{1}|C(\Omega_{2})}\bigg{)},$
then all Dirichlet eigenvalues of $a\Omega_{1}\times\Omega_{2}$ satisfy Pólya’s conjecture (2.1). This complete the proof of Theorem 1.1, part (1), for the case of $d_{2}\geq 3$.
3.2. The Dirichlet case with $d_{2}=2$
Since $C_{2}=\frac{1}{4\pi}$, the inequality (3.1) becomes
(3.5) $\sum_{l=1}^{Z^{\lambda}_{a}}\mathcal{N}^{D}_{\Omega_{2}}(\lambda-a^{-2}\lambda_{l}(\Omega_{1}))$
$\leq$ $(4\pi)^{-1}|\Omega_{2}|a^{-2}\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))+C(\Omega_{2})a^{-1}\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))^{\frac{1}{2}}$
and the inequality (3.2) gives, for $a^{2}\lambda>C(\Omega_{1})$,
(3.6) $\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))\leq L_{1,d_{1}}|\Omega_{1}|a^{d_{1}+2}\lambda^{\frac{d_{1}+2}{2}}-\frac{1}{5}L_{1,d_{1}-1}|\partial\Omega_{1}|a^{d_{1}+1}\lambda^{\frac{d_{1}+1}{2}}.$
To estimate the second term in (3.5), we use Li-Yau’s lower bound (1.5), namely
$\lambda_{l}(\Omega_{1})\geq\frac{d_{1}}{d_{1}+2}l^{\frac{2}{d_{1}}}(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}},$
to get
(3.7) $\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega_{1}))^{\frac{1}{2}}\leq$ $\sum_{l=1}^{Z^{\lambda}_{a}}\big{(}a^{2}\lambda-\frac{d_{1}}{d_{1}+2}l^{\frac{2}{d_{1}}}(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}}\big{)}^{\frac{1}{2}}$
$\leq$ $\int_{0}^{\infty}\big{(}a^{2}\lambda-\frac{d_{1}}{d_{1}+2}x^{\frac{2}{d_{1}}}(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}}\big{)}_{+}^{\frac{1}{2}}\mathrm{d}x$
$=$ $(\frac{d_{1}+2}{d_{1}})^{\frac{d_{1}}{2}}C_{d_{1}}|\Omega_{1}|\frac{d_{1}}{2}a^{d_{1}+1}\lambda^{\frac{d_{1}+1}{2}}\int_{0}^{1}(1-s)^{\frac{1}{2}}s^{\frac{d_{1}}{2}-1}\mathrm{d}s$
$=$ $(\frac{d_{1}+2}{d_{1}})^{\frac{d_{1}}{2}}L_{\frac{1}{2},d_{1}}|\Omega_{1}|a^{d_{1}+1}\lambda^{\frac{d_{1}+1}{2}}.$
Then by (3.5), (3.6) and (3.7), one has that if $a^{2}\lambda>C(\Omega_{1})$, then
$\sum_{l=1}^{Z^{\lambda}_{a}}\mathcal{N}^{D}_{\Omega_{2}}(\lambda-a^{-2}\lambda_{l}(\Omega_{1}))$
$\leq$ $C_{d_{1}+2}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+2}{2}}-\frac{1}{5}C_{d_{1}+1}|\Omega_{2}||\partial\Omega_{1}|a^{d_{1}-1}\lambda^{\frac{d_{1}+1}{2}}+$
$(\frac{d_{1}+2}{d_{1}})^{\frac{d_{1}}{2}}C(\Omega_{2})L_{\frac{1}{2},d_{1}}|\Omega_{1}|a^{d_{1}}\lambda^{\frac{d_{1}+1}{2}}.$
So if we assume
$a<\frac{C_{d_{1}+1}|\Omega_{2}||\partial\Omega_{1}|}{5(\frac{d_{1}+2}{d_{1}})^{\frac{d_{1}}{2}}C(\Omega_{2})L_{\frac{1}{2},d_{1}}|\Omega_{1}|}=\frac{1}{5\pi}\big{(}\frac{d_{1}}{d_{1}+2}\big{)}^{\frac{d_{1}}{2}}\frac{|\Omega_{2}||\partial\Omega_{1}|}{C(\Omega_{2})|\Omega_{1}|},$
then for any $\lambda>a^{-2}C(\Omega_{1})$, one also gets the demanded inequality
$\mathcal{N}^{D}_{a\Omega_{1}\times\Omega_{2}}(\lambda)\leq C_{d_{1}+2}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+2}{2}}.$
For $\lambda<a^{-2}C(\Omega_{1})$, one just repeat the corresponding part of the proof of the Dirichlet case with $d_{2}\geq 3$, so we omit it. This completes the proof of Theorem 1.1, (1).
3.3. The Neumann case
Since the Neumann eigenvalues of $a\Omega_{1}\times\Omega_{2}$ are
$a^{-2}\mu_{l}(\Omega_{1})+\mu_{k}(\Omega_{2}),\qquad\forall l,k\in\mathbb{Z}_{\geq 0},$
one has
$\mathcal{N}^{N}_{a\Omega_{1}\times\Omega_{2}}(\lambda)=\sum_{l=0}^{Y_{a}^{\lambda}}\mathcal{N}^{N}_{\Omega_{2}}(\lambda-a^{-2}\mu_{l}(\Omega_{1}))$
where
$Y_{a}^{\lambda}=\mathcal{N}^{N}_{\Omega_{1}}(a^{2}\lambda)-1.$
By inequality (2.7), there exists a constant $C_{1}(\Omega_{2})>0$ such that
$\mathcal{N}^{N}_{\Omega_{2}}(\lambda)\geq C_{d_{2}}|\Omega_{2}|\lambda^{\frac{d_{2}}{2}}-C_{1}(\Omega_{2})\lambda^{\frac{d_{2}-1}{2}},\qquad\forall\lambda>0.$
So we get
(3.8) $\sum_{l=0}^{Y_{a}^{\lambda}}\mathcal{N}^{N}_{\Omega_{2}}(\lambda-a^{-2}\mu_{l}(\Omega_{1}))$
$\geq$ $C_{d_{2}}|\Omega_{2}|\sum_{l=0}^{Y_{a}^{\lambda}}(\lambda-a^{-2}\mu_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}-C_{1}(\Omega_{2})\sum_{l=0}^{Y_{a}^{\lambda}}(\lambda-a^{-2}\mu_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}$
$=$ $C_{d_{2}}|\Omega_{2}|a^{-d_{2}}\sum_{l=0}^{Y_{a}^{\lambda}}(a^{2}\lambda-\mu_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}-C_{1}(\Omega_{2})a^{1-d_{2}}\sum_{l=0}^{Y_{a}^{\lambda}}(a^{2}\lambda-\mu_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}.$
For the first term, by (2.9), there exists a constant $C_{1}(\Omega_{1})>0$ such that if $a^{2}\lambda>C_{1}(\Omega_{1})$, then
(3.9) $\sum_{l=0}^{Y_{a}^{\lambda}}(a^{2}\lambda-\mu_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}\geq L_{\frac{d_{2}}{2},d_{1}}|\Omega_{1}|a^{d_{1}+d_{2}}\lambda^{\frac{d_{1}+d_{2}}{2}}+\frac{1}{5}L_{\frac{d_{2}}{2},d_{1}-1}|\partial\Omega_{1}|a^{d_{1}+d_{2}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
To estimate the second term, we use (2.5) to get $L=L(\Omega_{1})>0$ such that
$\mu_{l}(\Omega_{1})\geq\frac{1}{2}l^{\frac{2}{d_{1}}}(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}},\text{ if }l\geq L.$
Note that if $a^{2}\lambda$ is large enough, one has
$\sum_{l=L}^{3L-1}\big{(}a^{2}\lambda-\frac{1}{2}l^{\frac{2}{d_{1}}}(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}}\big{)}^{\frac{d_{2}-1}{2}}\geq L(a^{2}\lambda)^{\frac{d_{2}-1}{2}}\geq\sum_{l=0}^{L-1}(a^{2}\lambda-\mu_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}.$
So there exists a constant $C_{2}(\Omega_{1})>0$ such that if $a^{2}\lambda>C_{2}(\Omega_{1})$, then
(3.10) $\sum_{l=0}^{Y_{a}^{\lambda}}(a^{2}\lambda-\mu_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}$
$\leq$ $2\sum_{l=0}^{Y_{a}^{\lambda}}\big{(}a^{2}\lambda-\frac{1}{2}l^{\frac{2}{d_{1}}}(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}}\big{)}^{\frac{d_{2}-1}{2}}$
$=$ $2(a^{2}\lambda)^{\frac{d_{2}-1}{2}}+2\sum_{l=1}^{Y_{a}^{\lambda}}\big{(}a^{2}\lambda-\frac{1}{2}l^{\frac{2}{d_{1}}}(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}}\big{)}^{\frac{d_{2}-1}{2}}$
$\leq$ $4\int_{0}^{\infty}\big{(}a^{2}\lambda-\frac{1}{2}x^{\frac{2}{d_{1}}}(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}}\big{)}_{+}^{\frac{d_{2}-1}{2}}\mathrm{d}x$
$=$ $2^{\frac{d_{1}}{2}+2}C_{d_{1}}|\Omega_{1}|\frac{d_{1}}{2}a^{d_{1}+d_{2}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}\int_{0}^{1}(1-s)^{\frac{d_{2}-1}{2}}s^{\frac{d_{1}}{2}-1}\mathrm{d}s$
$=$ $2^{\frac{d_{1}}{2}+1}C_{d_{1}}B(\frac{d_{1}}{2},\frac{d_{2}+1}{2})|\Omega_{1}|d_{1}a^{d_{1}+d_{2}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
Thus by (3.8), (3.9) and (3.10), one has that if $a^{2}\lambda>\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))$, then
$\sum_{l=0}^{Y^{\lambda}_{a}}\mathcal{N}^{N}_{\Omega_{2}}(\lambda-a^{-2}\mu_{l}(\Omega_{1}))$
$\geq$ $C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}+\frac{1}{5}C_{d_{1}+d_{2}-1}|\Omega_{2}||\partial\Omega_{1}|a^{d_{1}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}-$
$C_{1}(\Omega_{2})2^{\frac{d_{1}}{2}+1}C_{d_{1}}B(\frac{d_{1}}{2},\frac{d_{2}+1}{2})|\Omega_{1}|d_{1}a^{d_{1}}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
So if we require
$a<\frac{C_{d_{1}+d_{2}-1}|\Omega_{2}||\partial\Omega_{1}|}{5C_{1}(\Omega_{2})2^{\frac{d_{1}}{2}+1}C_{d_{1}}B(\frac{d_{1}}{2},\frac{d_{2}+1}{2})|\Omega_{1}|d_{1}}=\frac{C_{d_{2}-1}|\Omega_{2}||\partial\Omega_{1}|}{5\cdot 2^{\frac{d_{1}}{2}+2}C_{1}(\Omega_{2})|\Omega_{1}|},$
then for any $\lambda>a^{-2}\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))$, one gets the demanded
$\mathcal{N}^{N}_{a\Omega_{1}\times\Omega_{2}}(\lambda)\geq C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}.$
Next, we consider $0<\lambda<a^{-2}\mu_{1}(\Omega_{1})$, in which case
$\mathcal{N}^{N}_{a\Omega_{1}\times\Omega_{2}}(\lambda)=\mathcal{N}^{N}_{\Omega_{2}}(\lambda).$
By Szegö-Weinberger inequality ([33], [34]), one has
$\mu_{1}(\Omega_{1})\leq(C_{d_{1}}|\Omega_{1}|)^{-\frac{2}{d_{1}}}$
which implies that for $0<\lambda<a^{-2}\mu_{1}(\Omega_{1})$,
$C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}$ $<C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|(a^{-2}\mu_{1}(\Omega_{1}))^{\frac{d_{1}}{2}}\lambda^{\frac{d_{2}}{2}}$
$\leq\frac{C_{d_{1}+d_{2}}}{C_{d_{1}}}|\Omega_{2}|\lambda^{\frac{d_{2}}{2}}.$
On the other hand, by (2.4) one has $\frac{C_{d_{1}+d_{2}}}{C_{d_{1}}}<C_{d_{2}}$. So by (2.5), there exists a constant $C_{2}(\Omega_{2})>0$ such that for $\lambda>C_{2}(\Omega_{2})$,
$\mathcal{N}^{N}_{\Omega_{2}}(\lambda)>\frac{C_{d_{1}+d_{2}}}{C_{d_{1}}}|\Omega_{2}|\lambda^{\frac{d_{2}}{2}}.$
Thus if $a^{-2}\mu_{1}(\Omega_{1})>\lambda>C_{2}(\Omega_{2})$, one gets
$\mathcal{N}^{N}_{a\Omega_{1}\times\Omega_{2}}(\lambda)=\mathcal{N}^{N}_{\Omega_{2}}(\lambda)>\frac{C_{d_{1}+d_{2}}}{C_{d_{1}}}|\Omega_{2}|\lambda^{\frac{d_{2}}{2}}\geq C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}.$
For $0<\lambda\leq C_{2}(\Omega_{2})$, we only need to require
$a<\big{(}C_{d_{1}+d_{2}}|\Omega_{1}||\Omega_{2}|C_{2}(\Omega_{2})^{\frac{d_{1}+d_{2}}{2}}\big{)}^{-\frac{1}{d_{1}}}=:C(\Omega_{1},\Omega_{2}),$
to get
$C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}<1\leq\mathcal{N}^{N}_{a\Omega_{1}\times\Omega_{2}}(\lambda).$
It remains to consider the case $a^{-2}\mu_{1}(\Omega_{1})\leq\lambda\leq a^{-2}\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))$ assuming $\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))>\mu_{1}(\Omega_{1})$. Let $\mu=a^{2}\lambda$, then by Corollary 2.4, one has
(3.11) $K_{1}(\Omega_{1})=\inf_{\mu_{1}(\Omega_{1})\leq\mu\leq\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))}\frac{\underset{\mu_{l}(\Omega_{1})<\mu}{\sum}(\mu-\mu_{l}(\Omega_{1}))^{\frac{d_{2}}{2}}-L_{\frac{d_{2}}{2},d_{1}}|\Omega_{1}|\mu^{\frac{d_{1}+d_{2}}{2}}}{\mu^{\frac{d_{1}+d_{2}-1}{2}}}>0.$
Let
(3.12) $K_{2}(\Omega_{1}):=\sup_{\mu_{1}(\Omega_{1})\leq\mu\leq\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))}\frac{\underset{\mu_{l}(\Omega_{1})<\mu}{\sum}(\mu-\mu_{l}(\Omega_{1}))^{\frac{d_{2}-1}{2}}}{\mu^{\frac{d_{1}+d_{2}-1}{2}}}>0.$
Then by (3.8) (3.11) and (3.12), for $a^{-2}\mu_{1}(\Omega_{1})\leq\lambda\leq a^{-2}\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))$ one has
$\sum_{l=0}^{Y_{a}^{\lambda}}\mathcal{N}^{N}_{\Omega_{2}}(\lambda-a^{-2}\mu_{l}(\Omega_{1}))$
$\geq$ $C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}+C_{d_{2}}|\Omega_{2}|K_{1}(\Omega_{1})a^{d_{1}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}-$
$C_{1}(\Omega_{2})K_{2}(\Omega_{1})a^{d_{1}}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
Thus if we assume
$a<\frac{C_{d_{2}}|\Omega_{2}|K_{1}(\Omega_{1})}{C_{1}(\Omega_{2})K_{2}(\Omega_{1})}$
then for any $a^{-2}\mu_{1}(\Omega_{1})\leq\lambda\leq a^{-2}\max(C_{1}(\Omega_{1}),C_{2}(\Omega_{1}))$, one gets the demanded inequality
$\mathcal{N}^{N}_{a\Omega_{1}\times\Omega_{2}}(\lambda)\geq C_{d_{1}+d_{2}}a^{d_{1}}|\Omega_{1}||\Omega_{2}|\lambda^{\frac{d_{1}+d_{2}}{2}}.$
Thus we conclude that for
$a<\min\bigg{(}\frac{C_{d_{2}-1}|\Omega_{2}||\partial\Omega_{1}|}{5\cdot 2^{\frac{d_{1}}{2}+2}C_{1}(\Omega_{2})|\Omega_{1}|},\frac{C_{d_{2}}|\Omega_{2}|K_{1}(\Omega_{1})}{C_{1}(\Omega_{2})K_{2}(\Omega_{1})},C(\Omega_{1},\Omega_{2})\bigg{)},$
[…]
4. Proof of Theorem 1.2
4.1. Two elementary lemmas
Before proving Theorem 1.2, we give two elementary lemmas that will play important roles later.
Lemma 4.1.
Let $f_{d}(x)=(\lambda-\frac{x^{2}\pi^{2}}{a^{2}})^{\frac{d}{2}}$, then
  1. $\mathrm{(1)}$
    $f_{d}$ is decreasing on $(0,\frac{a\sqrt{\lambda}}{\pi})$.
  2. $\mathrm{(2)}$
    If $d\geq 3$, $f_{d}$ is concave on $(0,\sqrt{\frac{\lambda}{d-1}}\frac{a}{\pi})$ and is convex on $(\sqrt{\frac{\lambda}{d-1}}\frac{a}{\pi},\frac{a\sqrt{\lambda}}{\pi})$.
  3. $\mathrm{(3)}$
    $\int_{0}^{\frac{a\sqrt{\lambda}}{\pi}}f_{d}(x)\mathrm{d}x=a\cdot\frac{C_{d+1}}{C_{d}}\lambda^{\frac{d+1}{2}}$.
Proof.
(1) is trivial. (2) follows from
$f_{d}^{\prime\prime}(x)=\frac{\pi^{2}d}{a^{2}}(\lambda-\frac{x^{2}\pi^{2}}{a^{2}})^{\frac{d}{2}-2}((d-1)\frac{\pi^{2}x^{2}}{a^{2}}-\lambda),$
and (3) is also elementary:
$\int_{0}^{\frac{a\sqrt{\lambda}}{\pi}}f_{d}(x)\mathrm{d}x=\lambda^{\frac{d+1}{2}}\frac{a}{\pi}\int_{0}^{1}(1-t^{2})^{\frac{d}{2}}\mathrm{d}t$
$=$ $\lambda^{\frac{d+1}{2}}\frac{a}{\pi}\int_{0}^{\frac{\pi}{2}}(\cos\theta)^{d+1}\mathrm{d}\theta=\lambda^{\frac{d+1}{2}}\frac{a}{\pi}\cdot\frac{\Gamma(\frac{1}{2})\Gamma(\frac{d}{2}+1)}{2\Gamma(\frac{d+1}{2}+1)}=a\cdot\frac{C_{d+1}}{C_{d}}\lambda^{\frac{d+1}{2}}.$
The second lemma is
Lemma 4.2.
Let
(4.1) $M_{a}^{\lambda}=\lfloor{\frac{a\sqrt{\lambda}}{\pi}}\rfloor,$
then for $\lambda\geq\frac{\pi^{2}}{a^{2}}$ (i.e. $M_{a}^{\lambda}\geq 1$), we have
(4.2) $\sum_{l=1}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})\leq\frac{2a\lambda^{\frac{3}{2}}}{3\pi}-\frac{\lambda}{8}-\frac{\sqrt{\lambda}\pi}{12a}$
and
(4.3) $\sum_{l=0}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})\geq\frac{2a}{3\pi}\lambda^{\frac{3}{2}}+\frac{1}{12}\lambda.$
Proof.
We have
$\sum_{l=1}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})=\lambda M_{a}^{\lambda}-\frac{\pi^{2}}{3a^{2}}(M_{a}^{\lambda})^{3}-\frac{\pi^{2}}{2a^{2}}(M_{a}^{\lambda})^{2}-\frac{\pi^{2}}{6a^{2}}M_{a}^{\lambda}.$
Let
(4.4) $g(x)=\lambda x-\frac{\pi^{2}}{3a^{2}}x^{3},$
then $g^{\prime}(x)=\lambda-\frac{\pi^{2}}{a^{2}}x^{2}$ which is positive if $x\in(0,\frac{a\sqrt{\lambda}}{\pi})$. So
$g(M^{\lambda}_{a})\leq g(\frac{a\sqrt{\lambda}}{\pi})=\frac{2a\lambda^{\frac{3}{2}}}{3\pi}.$
Combining with the fact
$\lfloor{x}\rfloor\geq\frac{x}{2},\qquad\forall x\geq 1,$
one gets (4.2). Similarly,
$\sum_{l=0}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})=\lambda+\lambda M_{a}^{\lambda}-\frac{\pi^{2}}{3a^{2}}(M_{a}^{\lambda})^{3}-\frac{\pi^{2}}{a^{2}}\cdot\frac{3(M_{a}^{\lambda})^{2}+M_{a}^{\lambda}}{6}.$
Again consider the function $g(x)$ defined in (4.4). Since $g^{\prime}(x)$ is positive and monotonically decreasing on $(0,\frac{a\sqrt{\lambda}}{\pi})$, one has
$\frac{2a\lambda^{\frac{3}{2}}}{3\pi}-(\lambda M_{a}^{\lambda}-\frac{\pi^{2}}{3a^{2}}(M_{a}^{\lambda})^{3})=g(\frac{a\sqrt{\lambda}}{\pi})-g(M_{a}^{\lambda})\leq g^{\prime}(M_{a}^{\lambda})=\lambda-\frac{\pi^{2}}{a^{2}}(M_{a}^{\lambda})^{2},$
which implies (4.3). ∎
Now we start to prove Theorem 1.2. Again by [12, Proposition 3.1], it is enough to assume that $\Omega$ is connected. Since any rectangle in $\mathbb{R}^{2}$ satisfies Pólya’s conjecture, we can assume that the dimension $d$ of $\Omega$ is at least 2. Again we argue by treating $d\geq 3$ and $d=2$ separately, and by treating Dirichlet case and Neumann case separately.
4.2. The Dirichlet case with $d=2$
The Dirichlet eigenvalues of $(0,a)\times\Omega$ are
$\frac{l^{2}\pi^{2}}{a^{2}}+\lambda_{k}(\Omega),\qquad l,k\in\mathbb{Z}_{>0},$
and thus
(4.5) $\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)=\sum_{l=1}^{M_{a}^{\lambda}}\mathcal{N}^{D}_{\Omega}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}}).$
Note that if $0<\lambda<\frac{\pi^{2}}{a^{2}}$, then
$\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)=0<C_{3}a|\Omega|\lambda^{\frac{3}{2}}.$
So one only need to consider the case $\lambda\geq\frac{\pi^{2}}{a^{2}}$, i.e. $M_{a}^{\lambda}\geq 1$. By inequality (2.6), for any $\lambda>0$, there exists a constant $C(\Omega)>0$ such that
$\mathcal{N}^{D}_{\Omega}(\lambda)\leq\frac{|\Omega|}{4\pi}\lambda+C(\Omega)\lambda^{\frac{1}{2}}.$
In view of (4.2) and the fact
$\sum_{l=1}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})^{\frac{1}{2}}\leq\int_{0}^{\frac{a\sqrt{\lambda}}{\pi}}(\lambda-\frac{x^{2}\pi^{2}}{a^{2}})^{\frac{1}{2}}\mathrm{d}x=\frac{a}{4}\lambda$
we get
$\sum_{l=1}^{M_{a}^{\lambda}}\mathcal{N}^{D}_{\Omega}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})\leq$ $\frac{|\Omega|}{4\pi}\sum_{l=1}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})+C(\Omega)\sum_{l=1}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})^{\frac{1}{2}}$
$\leq$ $\frac{a|\Omega|\lambda^{\frac{3}{2}}}{6\pi^{2}}-\frac{|\Omega|\lambda}{32\pi}+\frac{C(\Omega)a}{4}\lambda.$
Thus if we assume $a<\frac{|\Omega|}{8\pi C(\Omega)}$, then
$\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)\leq\frac{a|\Omega|\lambda^{\frac{3}{2}}}{6\pi^{2}}=C_{3}a|\Omega|\lambda^{\frac{3}{2}}.$
This completes the proof of the Dirichlet case with $d=2$.
4.3. The Neumann case with $d=2$
For the Neumann case, the eigenvalues of $(0,a)\times\Omega$ are
$\frac{l^{2}\pi^{2}}{a^{2}}+\mu_{k}(\Omega),\qquad l,k\in\mathbb{Z}_{\geq 0},$
thus
(4.6) $\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)=\sum_{l=0}^{M_{a}^{\lambda}}\mathcal{N}^{N}_{\Omega}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}}).$
By inequality (2.6), for any $\lambda>0$, there exists $C(\Omega)>0$ such that
$\mathcal{N}^{N}_{\Omega}(\lambda)\geq C_{d}|\Omega|\lambda-C(\Omega)\lambda^{\frac{1}{2}}.$
In view of (4.3) and the fact
$\sum_{l=0}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})^{\frac{1}{2}}\leq\lambda^{\frac{1}{2}}+\int_{0}^{\frac{a\sqrt{\lambda}}{\pi}}(\lambda-\frac{x^{2}\pi^{2}}{a^{2}})^{\frac{1}{2}}\mathrm{d}x=\lambda^{\frac{1}{2}}+\frac{a}{4}\lambda$
we get, for $\lambda\geq\frac{\pi^{2}}{a^{2}}$,
$\sum_{l=0}^{M_{a}^{\lambda}}\mathcal{N}^{N}_{\Omega}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})\geq$ $\frac{|\Omega|}{4\pi}\sum_{l=0}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})-C(\Omega)\sum_{l=0}^{M_{a}^{\lambda}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})^{\frac{1}{2}}$
$\geq$ $\frac{a|\Omega|\lambda^{\frac{3}{2}}}{6\pi^{2}}+\frac{|\Omega|\lambda}{48\pi}-C(\Omega)(\lambda^{\frac{1}{2}}+\frac{a\lambda}{4}).$
So if we assume $a\leq\frac{|\Omega|}{96C(\Omega)}$, then $\lambda\geq\frac{\pi^{2}}{a^{2}}\geq(\frac{96\pi C(\Omega)}{|\Omega|})^{2}$ and thus
$\frac{C(\Omega)a\lambda}{4}\leq\frac{|\Omega|\lambda}{4\cdot 96}<\frac{|\Omega|\lambda}{96\pi}\quad\text{and}\quad C(\Omega)\lambda^{\frac{1}{2}}\leq\frac{|\Omega|\lambda}{96\pi}.$
In other words, if we assume $a\leq\frac{|\Omega|}{96C(\Omega)}$, then for any $\lambda\geq\frac{\pi^{2}}{a^{2}}$,
$\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)\geq\frac{a|\Omega|\lambda^{\frac{3}{2}}}{6\pi^{2}}=C_{3}a|\Omega|\lambda^{\frac{3}{2}}.$
Next, if $0<\lambda<\frac{\pi^{2}}{a^{2}}$, then
$\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)=\mathcal{N}^{N}_{\Omega}(\lambda),\quad\text{ and }\quad C_{3}a|\Omega|\lambda^{\frac{3}{2}}<C_{3}\pi|\Omega|\lambda.$
By (2.4), one has $C_{3}\pi<C_{2}$. So by (2.5), there exists a constant $C_{1}(\Omega)>0$, such that if $\lambda\geq C_{1}(\Omega)$, then
$\mathcal{N}^{N}_{\Omega}(\lambda)>C_{3}\pi|\Omega|\lambda.$
Thus for $C_{1}(\Omega)\leq\lambda<\frac{\pi^{2}}{a^{2}}$, one gets
$\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)=\mathcal{N}^{N}_{\Omega}(\lambda)>C_{3}\pi|\Omega|\lambda>C_{3}a|\Omega|\lambda^{\frac{3}{2}}.$
Finally for $0<\lambda\leq C_{1}(\Omega)$, we may require $a\leq(C_{3}|\Omega|)^{-1}C_{1}(\Omega)^{-\frac{3}{2}}$ to get
$C_{3}a|\Omega|\lambda^{\frac{3}{2}}\leq 1\leq\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda).$
Combining all discussions above, one has that if
(4.7) $a<\min\bigg{(}\frac{|\Omega|}{96C(\Omega)},(C_{3}|\Omega|)^{-1}C_{1}(\Omega)^{-\frac{3}{2}}\bigg{)},$
then all Neumann eigenvalues of $(0,a)\times\Omega$ satisfy Pólya’s conjecture (1.3). Thus we complete the proof of Theorem 1.2 with $d=2$.
4.4. The Dirichlet case with $d\geq 3$
Again we have
$\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)=\sum_{l=1}^{M_{a}^{\lambda}}\mathcal{N}^{D}_{\Omega}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})$
and there exists a constant $C(\Omega)>0$ such that
$\mathcal{N}^{D}_{\Omega}(\lambda)\leq C_{d}|\Omega|\lambda^{\frac{d}{2}}+C(\Omega)\lambda^{\frac{d-1}{2}}.$
So we get
(4.8) $\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)\leq C_{d}|\Omega|\sum_{l=1}^{M_{a}^{\lambda}}f_{d}(l)+C(\Omega)\sum_{l=1}^{M_{a}^{\lambda}}f_{d-1}(l),$
where $f_{d}$ is defined in Lemma 4.1. We split the first sum into two parts. Denote
$N_{a}^{\lambda}=\lfloor\sqrt{\frac{\lambda}{d-1}}\frac{a}{\pi}\rfloor.$
By concavity of $f_{d}$ (see (2) of Lemma 4.1), one has
$\int_{0}^{N_{a}^{\lambda}}f_{d}(x)\mathrm{d}x-\sum_{l=1}^{N_{a}^{\lambda}}f_{d}(l)\geq\sum_{l=0}^{N_{a}^{\lambda}-1}f_{d}(l)-\int_{0}^{N_{a}^{\lambda}}f_{d}(x)\mathrm{d}x$
which implies
$\sum_{l=1}^{N_{a}^{\lambda}}f_{d}(l)\leq\int_{0}^{N_{a}^{\lambda}}f_{d}(x)\mathrm{d}x-\frac{1}{2}\bigg{(}\lambda^{\frac{d}{2}}-f_{d}(N_{a}^{\lambda})\bigg{)}.$
First consider $\lambda\geq\frac{(d-1)\pi^{2}}{a^{2}}$, in which case $N_{a}^{\lambda}\geq\frac{1}{2}\frac{a}{\pi}\sqrt{\frac{\lambda}{d-1}}$, and thus
$f_{d}(N_{a}^{\lambda})\leq f_{d}(\frac{1}{2}\frac{a}{\pi}\sqrt{\frac{\lambda}{d-1}})=(\frac{4d-5}{4d-4})^{\frac{d}{2}}\lambda^{\frac{d}{2}}.$
For simplicity, we denote
$A_{d}=\frac{1}{2}\bigg{(}1-\big{(}\frac{4d-5}{4d-4}\big{)}^{\frac{d}{2}}\bigg{)}C_{d}|\Omega|.$
Then we get, for $\lambda\geq\frac{(d-1)\pi^{2}}{a^{2}}$,
$\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)\leq$ $C_{d}|\Omega|\sum_{l=1}^{N_{a}^{\lambda}}f_{d}(l)+C_{d}|\Omega|\sum_{l=N_{a}^{\lambda}+1}^{M_{a}^{\lambda}}f_{d}(l)+C(\Omega)\sum_{l=1}^{M_{a}^{\lambda}}f_{d-1}(l)$
$\leq$ $C_{d}|\Omega|\int_{0}^{{\frac{a\sqrt{\lambda}}{\pi}}}f_{d}(x)\mathrm{d}x-A_{d}\lambda^{\frac{d}{2}}+C(\Omega)\int_{0}^{{\frac{a\sqrt{\lambda}}{\pi}}}f_{d-1}(x)\mathrm{d}x$
$=$ $C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}}-A_{d}\lambda^{\frac{d}{2}}+C(\Omega)\frac{C_{d}a}{C_{d-1}}\lambda^{\frac{d}{2}}.$
So if we assume $a\leq\frac{A_{d}\cdot C_{d-1}}{C(\Omega)\cdot C_{d}}$, then for any $\lambda\geq\frac{(d-1)\pi^{2}}{a^{2}_{1}}$,
(4.9) $\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)\leq C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}}.$
Note that if $0<\lambda<\frac{\pi^{2}}{a^{2}}$, then we automatically have
$\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)=0<C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}},$
so it remains to consider the case $\frac{\pi^{2}}{a^{2}}\leq\lambda<\frac{(d-1)\pi^{2}}{a^{2}}$. Let $\mu=\frac{\lambda a^{2}}{\pi^{2}}$, then $11\leq\mu<d-$ 1. Let
$H_{1}:=\inf_{1\leq\mu<d-1}\frac{\int^{\sqrt{\mu}}_{0}(\mu-x^{2})^{\frac{d}{2}}\mathrm{d}x-\sum_{0<l^{2}<\mu}(\mu-l^{2})^{\frac{d}{2}}}{\mu^{\frac{d}{2}}}>0,$
then
$\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)\leq$ $C_{d}|\Omega|\int_{0}^{{\frac{a\sqrt{\lambda}}{\pi}}}f_{d}(x)\mathrm{d}x-C_{d}|\Omega|H_{1}\lambda^{\frac{d}{2}}+C(\Omega)\int_{0}^{{\frac{a\sqrt{\lambda}}{\pi}}}f_{d-1}(x)\mathrm{d}x$
$=$ $C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}}-C_{d}|\Omega|H_{1}\lambda^{\frac{d}{2}}+C(\Omega)\frac{C_{d}a}{C_{d-1}}\lambda^{\frac{d}{2}}.$
So if we assume $a\leq\frac{C_{d-1}|\Omega|H_{1}}{C(\Omega)}$, then for any $\frac{\pi^{2}}{a^{2}}\leq\lambda<\frac{(d-1)\pi^{2}}{a^{2}}$,
$\mathcal{N}^{D}_{(0,a)\times\Omega}(\lambda)\leq C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}}.$
Combining with (4.9), one gets that if
$a\leq\min\bigg{(}\frac{A_{d}\cdot C_{d-1}}{C(\Omega)\cdot C_{d}},\frac{C_{d-1}|\Omega|H_{1}}{C(\Omega)}\bigg{)},$
then all Dirichlet eigenvalues of $(0,a)\times\Omega$ satisfy Pólya’s conjecture (1.2).
4.5. The Neumann case with $d\geq 3$
Again we have
$\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)=\sum_{l=0}^{M_{a}^{\lambda}}\mathcal{N}^{N}_{\Omega}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})$
and there exists a constant $C(\Omega)>0$ such that
$\mathcal{N}^{N}_{\Omega}(\lambda)\geq C_{d}|\Omega|\lambda^{\frac{d}{2}}-C(\Omega)\lambda^{\frac{d-1}{2}},\qquad\forall\lambda>0.$
So we get
(4.10) $\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)\geq C_{d}|\Omega|\sum_{l=0}^{M_{a}^{\lambda}}f_{d}(l)-C(\Omega)\sum_{l=0}^{M_{a}^{\lambda}}f_{d-1}(l).$
By convexity of $f_{d}$ (see (2) of Lemma 4.1), one has
$\sum_{l=N_{a}^{\lambda}+1}^{M_{a}^{\lambda}}f_{d}(l)-\int_{N_{a}^{\lambda}+1}^{\frac{a\sqrt{\lambda}}{\pi}}f_{d}(x)\mathrm{d}x\geq\int_{N_{a}^{\lambda}+1}^{\frac{a\sqrt{\lambda}}{\pi}}f_{d}(x)\mathrm{d}x-\sum_{l=N_{a}^{\lambda}+2}^{M_{a}^{\lambda}}f_{d}(l)$
which implies
$\sum_{l=N_{a}^{\lambda}+1}^{M_{a}^{\lambda}}f_{d}(l)\geq\int_{N_{a}^{\lambda}+1}^{\frac{a\sqrt{\lambda}}{\pi}}f_{d}(x)\mathrm{d}x+\frac{1}{2}f_{d}(N_{a}^{\lambda}+1).$
If $\lambda\geq\frac{9\pi^{2}(d-1)}{a^{2}}$, then $N_{a}^{\lambda}\geq 3$ and thus $N_{a}^{\lambda}+1\leq\frac{4a}{3\pi}\sqrt{\frac{\lambda}{d-1}}$, which implies
$f_{d}(N_{a}^{\lambda}+1)\geq f_{d}(\frac{4a}{3\pi}\sqrt{\frac{\lambda}{d-1}})\geq 3^{-{d}}\lambda^{\frac{d}{2}},$
where we used $d\geq 3$. For simplicity, we denote
$B_{d}=\frac{1}{2}3^{-d}C_{d}|\Omega|.$
Then for $\lambda\geq\frac{9\pi^{2}(d-1)}{a^{2}}$,
$\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)\geq$ $C_{d}|\Omega|\sum_{l=0}^{N_{a}^{\lambda}}f_{d}(l)+C_{d}|\Omega|\sum_{l=N_{a}^{\lambda}+1}^{M_{a}^{\lambda}}f_{d}(l)-C(\Omega)\sum_{l=0}^{M_{a}^{\lambda}}f_{d-1}(l)$
$\geq$ $C_{d}|\Omega|\int_{0}^{\frac{a\sqrt{\lambda}}{\pi}}\!\!\!f_{d}(x)\mathrm{d}x+B_{d}\lambda^{\frac{d}{2}}-C(\Omega)\big{(}\lambda^{\frac{d-1}{2}}+\int_{0}^{{\frac{a\sqrt{\lambda}}{\pi}}}\!\!\!f_{d-1}(x)\mathrm{d}x\big{)}$
$=$ $C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}}+B_{d}\lambda^{\frac{d}{2}}-C(\Omega)\big{(}\lambda^{\frac{d-1}{2}}+\frac{C_{d}a}{C_{d-1}}\lambda^{\frac{d}{2}}\big{)}.$
So if we assume
$a\leq\min\bigg{(}\frac{B_{d}C_{d-1}}{2C(\Omega)C_{d}},\frac{B_{d}3\pi\sqrt{d-1}}{2C(\Omega)}\bigg{)},$
then $\lambda\geq\frac{9\pi^{2}(d-1)}{a^{2}}\geq\frac{4C(\Omega)^{2}}{B_{d}^{2}}$ and thus
$\frac{C(\Omega)C_{d}a}{C_{d-1}}\lambda^{\frac{d}{2}}\leq\frac{1}{2}B_{d}\lambda^{\frac{d}{2}}\quad\text{and}\quad C(\Omega)\lambda^{\frac{d-1}{2}}\leq\frac{1}{2}B_{d}\lambda^{\frac{d}{2}}.$
Thus for any $\lambda\geq\frac{9\pi^{2}(d-1)}{a^{2}}$, one gets the demanded inequality
$\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)\geq C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}}.$
Next if $\frac{\pi^{2}}{a^{2}}\leq\lambda\leq\frac{9\pi^{2}(d-1)}{a^{2}}$, let $\mu=\frac{\lambda a^{2}}{\pi^{2}}$, then $1\leq\mu\leq 9(d-1)$. Let
$H_{2}=\inf_{1\leq\mu\leq 9(d-1)}\frac{\sum_{0\leq l^{2}<\mu}(\mu-l^{2})^{\frac{d}{2}}-\int_{0}^{\sqrt{\mu}}(\mu-x^{2})^{\frac{d}{2}}\mathrm{d}x}{\mu^{\frac{d}{2}}}>0,$
then
$\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)\geq$ $C_{d}|\Omega|\int_{0}^{\frac{a\sqrt{\lambda}}{\pi}}\!\!\!f_{d}(x)\mathrm{d}x+C_{d}|\Omega|H_{2}\lambda^{\frac{d}{2}}-C(\Omega)\big{(}\lambda^{\frac{d-1}{2}}+\int_{0}^{{\frac{a\sqrt{\lambda}}{\pi}}}\!\!\!f_{d-1}(x)\mathrm{d}x\big{)}$
$=$ $C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}}+C_{d}|\Omega|H_{2}\lambda^{\frac{d}{2}}-C(\Omega)\big{(}\lambda^{\frac{d-1}{2}}+\frac{C_{d}a}{C_{d-1}}\lambda^{\frac{d}{2}}\big{)}.$
Similar to the case $\lambda\geq\frac{9\pi^{2}(d-1)}{a^{2}}$, one can prove that if we take $a\leq\min\big{(}\frac{|\Omega|H_{2}}{2C_{d-1}},\frac{\pi C_{d}|\Omega|H_{2}}{2C(\Omega)}\big{)}$, then for any $\frac{\pi^{2}}{a^{2}}\leq\lambda\leq\frac{9\pi^{2}(d-1)}{a^{2}}$, one gets
$\mathcal{N}^{N}_{(0,a)\times\Omega}(\lambda)\geq C_{d+1}a|\Omega|\lambda^{\frac{d+1}{2}}.$
Finally for $0<\lambda<\frac{\pi^{2}}{a^{2}}$, we just repeat the corresponding part in the proof of the Neumann case with $d=2$. In conclusion, we get: if
$a<\min\bigg{(}\frac{B_{d}C_{d-1}}{2C(\Omega)C_{d}},\frac{B_{d}3\pi\sqrt{d-1}}{2C(\Omega)},\frac{|\Omega|H_{2}}{2C_{d-1}},\frac{\pi C_{d}|\Omega|H_{2}}{2C(\Omega)},(C_{d+1}|\Omega|)^{-1}C_{1}(\Omega)^{-\frac{d+1}{2}}\bigg{)},$
[…]
5. Proof of Theorem 1.3
Again by [12, Proposition 3.1], it is enough to assume that both $\Omega$ and $M$ are connected. Let the eigenvalues of $M$ be
$0=\lambda_{0}(M)<\lambda_{1}(M)\leq\cdots\nearrow\infty,$
and the counting functions for the eigenvalues of $M$ be
$\mathcal{N}_{M}(\lambda)=\#\{n|\ \lambda_{n}(M)<\lambda\}.$
B. M. Levitan ([22]) and V. G. Avakumović ([1]) proved that
$\mathcal{N}_{M}(\lambda)=C_{d_{2}}|M|\lambda^{\frac{d_{2}}{2}}+\mathrm{O}(\lambda^{\frac{d-1}{2}}),\text{ as }\lambda\to\infty.$
So there exists a constant $C(M)>0$ such that
(5.1) $\mathcal{N}_{M}(\lambda)\geq C_{d_{2}}|M|\lambda^{\frac{d_{2}}{2}}-C(M)\lambda^{\frac{d_{2}-1}{2}},\ \forall\lambda>0.$
Repeating the proof of the Neumann case of Theorem 1.1 and Theorem 1.2 word by word, one can easily prove the Neumann case of Theorem 1.3.
For the Dirichlet case of Theorem 1.3, since 0 is an eigenvalue of $M$, one can only get that there exists a constant $C_{1}(M)>0$ such that
(5.2) $\mathcal{N}_{M}(\lambda)\leq C_{d_{2}}|M|\lambda^{\frac{d_{2}}{2}}+C_{1}(M)\lambda^{\frac{d_{2}-1}{2}}+1,\ \forall\lambda>0.$
So to prove the Dirichlet case of Theorem 1.3, one need to carefully handle this extra number. Again we divided the proof into three parts: the Dirichlet case with $d_{1}=1$ and $d_{2}=2$, the Dirichlet case with $d_{1}=1$ and $d_{2}\geq 3$, and the Dirichlet case with $d_{1}\geq 2$.
5.1. The Dirichlet case with $d_{1}=1$ and $d_{2}=2$
When $d_{1}=1$, we can assume $\Omega=(0,1)$ for simplicity. The Dirichlet eigenvalues of $(0,a)\times M$ are
$\frac{l^{2}\pi^{2}}{a^{2}}+\lambda_{k}(M),\qquad l\in\mathbb{Z}_{>0},\ k\in\mathbb{Z}_{\geq 0}.$
If $0<\lambda<\frac{\pi^{2}}{a^{2}}$, then $\mathcal{N}^{D}_{(0,a)\times M}(\lambda)=0<C_{3}a|M|\lambda^{\frac{3}{2}}$. For $\lambda\geq\frac{\pi^{2}}{a^{2}}$, by (4.2) we get
$\mathcal{N}^{D}_{(0,a)\times M}(\lambda)$ $=\sum_{l=1}^{M^{\lambda}_{a}}\mathcal{N}_{M}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})$
$\leq\frac{|M|}{4\pi}\sum_{l=1}^{M^{\lambda}_{a}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})+C_{1}(M)\sum_{l=1}^{M^{\lambda}_{a}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})^{\frac{1}{2}}+M^{\lambda}_{a}$
$\leq\frac{|M|}{4\pi}(\frac{2a\lambda^{\frac{3}{2}}}{3\pi}-\frac{\lambda}{8}-\frac{\sqrt{\lambda}\pi}{12a})+\frac{C_{1}(M)a}{4}\lambda+\frac{a\sqrt{\lambda}}{\pi}.$
Note that if $a\leq\sqrt{\frac{|M|\pi}{48}}$, then
$\frac{|M|}{4\pi}(\frac{2a\lambda^{\frac{3}{2}}}{3\pi}-\frac{\lambda}{8}-\frac{\sqrt{\lambda}\pi}{12a})+\frac{C_{1}(M)a}{4}\lambda+\frac{a\sqrt{\lambda}}{\pi}\leq\frac{a|M|\lambda^{\frac{3}{2}}}{6\pi^{2}}-\frac{|M|\lambda}{32\pi}+\frac{C_{1}(M)a}{4}\lambda.$
Thus we proved: if
(5.3) $a\leq\min\big{(}\sqrt{\frac{|M|\pi}{48}},\frac{|M|}{8\pi C_{1}(M)}\big{)},$
then all Dirichlet eigenvalues of $(0,a)\times M$ satisfy Pólya’s conjecture (1.2).
5.2. The Dirichlet case with $d_{1}=1$ and $d_{2}\geq 3$
We still have
$\mathcal{N}^{D}_{(0,a)\times M}(\lambda)$ $=\sum_{l=1}^{M^{\lambda}_{a}}\mathcal{N}_{M}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})$
$\leq C_{d_{2}}|M|\sum_{l=1}^{M^{\lambda}_{a}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})^{\frac{d_{2}}{2}}+C_{1}(M)\sum_{l=1}^{M^{\lambda}_{a}}(\lambda-\frac{l^{2}\pi^{2}}{a^{2}})^{\frac{d_{2}-1}{2}}+M^{\lambda}_{a}.$
As in §4.4, if $\lambda\geq\frac{(d_{2}-1)\pi^{2}}{a^{2}}$, one has
$\mathcal{N}^{D}_{(0,a)\times M}(\lambda)\leq C_{d_{2}+1}a|M|\lambda^{\frac{d_{2}+1}{2}}-A_{d_{2}}\lambda^{\frac{d_{2}}{2}}+C_{1}(M)\frac{C_{d_{2}}a}{C_{d_{2}-1}}\lambda^{\frac{d_{2}}{2}}+\frac{a\sqrt{\lambda}}{\pi},$
where $A_{d_{2}}=\frac{1}{2}(1-(\frac{4d_{2}-5}{4d_{2}-4})^{\frac{d_{2}}{2}})C_{d_{2}}|M|$. To control the last term, we require
$a<\big{(}\frac{\pi}{2}A_{d_{2}}\big{)}^{\frac{1}{d_{2}}}\big{(}\pi^{2}(d_{2}-1)\big{)}^{\frac{d_{2}-1}{2d_{2}}}$
to get
$\frac{a\sqrt{\lambda}}{\pi}\leq\frac{1}{2}A_{d_{2}}\lambda^{\frac{d_{2}}{2}}$
for all $\lambda\geq\frac{(d_{2}-1)\pi^{2}}{a^{2}}$. Repeating §4.4, we will get: if
$a<\min\bigg{(}\pi\big{(}\frac{1}{2}A_{d_{2}}\big{)}^{\frac{1}{d_{2}}}(d_{2}-1)^{\frac{d_{2}-1}{2d_{2}}},\frac{A_{d_{2}}C_{d_{2}-1}}{2C_{1}(M)C_{d_{2}}}\bigg{)},$
then for any $\lambda\geq\frac{(d_{2}-1)\pi^{2}}{a^{2}}$,
$\mathcal{N}^{D}_{(0,a)\times M}(\lambda)\leq C_{d_{2}+1}a|M|\lambda^{\frac{d_{2}+1}{2}}.$
For $\lambda<\frac{(d_{2}-1)\pi^{2}}{a^{2}}$, again one only needs to consider $\frac{\pi^{2}}{a^{2}}\leq\lambda<\frac{(d_{2}-1)\pi^{2}}{a^{2}}$. As in §4.4, in this case one has
$\mathcal{N}^{D}_{(0,a)\times M}(\lambda)\leq C_{d_{2}+1}a|M|\lambda^{\frac{d_{2}+1}{2}}-C_{d_{2}}|M|H_{1}\lambda^{\frac{d_{2}}{2}}+C_{1}(M)\frac{C_{d_{2}}a}{C_{d_{2}-1}}\lambda^{\frac{d_{2}}{2}}+\frac{a\sqrt{\lambda}}{\pi},$
where
$H_{1}:=\inf_{1\leq\mu\leq d_{2}-1}\frac{\int^{\sqrt{\mu}}_{0}(\mu-x^{2})^{\frac{d_{2}}{2}}\mathrm{d}x-\sum_{0<l^{2}<\mu}(\mu-l^{2})^{\frac{d_{2}}{2}}}{\mu^{\frac{d_{2}}{2}}}>0.$
So if we assume
$a<\min\bigg{(}\pi\big{(}\frac{1}{2}C_{d_{2}}|M|H_{1}\big{)}^{\frac{1}{d_{2}}},\frac{C_{d_{2}-1}|M|H_{1}}{2C_{1}(M)}\bigg{)},$
then for all $\frac{\pi^{2}}{a^{2}}\leq\lambda<\frac{(d_{2}-1)\pi^{2}}{a^{2}}$, one has
$\mathcal{N}^{D}_{(0,a)\times M}(\lambda)\leq C_{d_{2}+1}a|M|\lambda^{\frac{d_{2}+1}{2}}.$
Thus if $d_{2}\geq 3$ and
$a<\min\bigg{(}\pi\big{(}\frac{1}{2}A_{d_{2}}\big{)}^{\frac{1}{d_{2}}}(d_{2}-1)^{\frac{d_{2}-1}{2d_{2}}},\frac{A_{d_{2}}C_{d_{2}-1}}{2C_{1}(M)C_{d_{2}}},\pi\big{(}\frac{1}{2}C_{d_{2}}|M|H_{1}\big{)}^{\frac{1}{d_{2}}},\frac{C_{d_{2}-1}|M|H_{1}}{2C_{1}(M)}\bigg{)},$
then all Dirichlet eigenvalues of $(0,a)\times M$ satisfy Pólya’s Conjecture (1.2).
5.3. The Dirichlet case with $d_{1}\geq 2$
The Dirichlet eigenvalues of $a\Omega\times M$ are
$a^{-2}\lambda_{l}(\Omega)+\lambda_{k}(M),\qquad l\in\mathbb{Z}_{>0},\ k\in\mathbb{Z}_{\geq 0},$
which implies
$\mathcal{N}^{D}_{a\Omega\times M}(\lambda)\leq C_{d_{2}}|M|a^{-d_{2}}\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega))^{\frac{d_{2}}{2}}+C_{1}(M)a^{1-d_{2}}\sum_{l=1}^{Z^{\lambda}_{a}}(a^{2}\lambda-\lambda_{l}(\Omega))^{\frac{d_{2}-1}{2}}+Z^{\lambda}_{a}$
where $Z^{\lambda}_{a}=\mathcal{N}^{D}_{\Omega}(a^{2}\lambda)$. To control the extra $Z^{\lambda}_{a}$, we use Li-Yau’s estimate (1.5) to get
$\mathcal{N}^{D}_{\Omega}(\lambda)\leq\big{(}\frac{d_{1}+2}{d_{1}}\big{)}^{\frac{d_{1}}{2}}C_{d_{1}}|\Omega|\lambda^{\frac{d_{1}}{2}},\qquad\forall\lambda>0.$
Thus for all $a>0$ and $\lambda>0$,
$Z^{\lambda}_{a}=\mathcal{N}^{D}_{\Omega}(a^{2}\lambda)\leq\big{(}\frac{d_{1}+2}{d_{1}}\big{)}^{\frac{d_{1}}{2}}C_{d_{1}}|\Omega|a^{d_{1}}\lambda^{\frac{d_{1}}{2}}.$
If $d_{2}\geq 3$, then as in §3.1, there exists a constant $C(\Omega)>0$ such that for $a^{2}\lambda>C(\Omega)$,
$\mathcal{N}^{D}_{a\Omega\times M}(\lambda)\leq$ $C_{d_{1}+d_{2}}a^{d_{1}}|\Omega||M|\lambda^{\frac{d_{1}+d_{2}}{2}}-\frac{1}{5}C_{d_{1}+d_{2}-1}|\partial\Omega||M|a^{d_{1}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}+$
$L_{\frac{d_{2}-1}{2},d_{1}}|\Omega|C_{1}(M)a^{d_{1}}\lambda^{\frac{d_{1}+d_{2}-1}{2}}+(\frac{d_{1}+2}{d_{1}})^{\frac{d_{1}}{2}}C_{d_{1}}|\Omega|a^{d_{1}}\lambda^{\frac{d_{1}}{2}}.$
So if we assume
$a<\big{(}\frac{C_{d_{1}+d_{2}-1}|\partial\Omega||M|}{10C_{d_{1}}|\Omega|}\big{)}^{\frac{1}{d_{2}}}\big{(}\frac{d_{1}+2}{d_{1}}\big{)}^{-\frac{1}{2}}C(\Omega)^{\frac{d_{2}-1}{2d_{2}}},$
then for any $\lambda>a^{-2}C(\Omega)$, the extra term is controlled by
$(\frac{d_{1}+2}{d_{1}})^{\frac{d_{1}}{2}}C_{d_{1}}|\Omega|a^{d_{1}}\lambda^{\frac{d_{1}}{2}}\leq\frac{1}{10}C_{d_{1}+d_{2}-1}|\partial\Omega||M|a^{d_{1}-1}\lambda^{\frac{d_{1}+d_{2}-1}{2}}.$
If $d_{2}=2$, then as in §3.2, there exists a constant $C(\Omega)>0$ such that if $a^{2}\lambda>C(\Omega)$, one has
$\mathcal{N}^{D}_{a\Omega\times M}(\lambda)\leq$ $C_{d_{1}+2}a^{d_{1}}|\Omega||M|\lambda^{\frac{d_{1}+2}{2}}-\frac{1}{5}C_{d_{1}+1}|M||\partial\Omega|a^{d_{1}-1}\lambda^{\frac{d_{1}+1}{2}}+$
$\big{(}\frac{d_{1}+2}{d_{1}}\big{)}^{\frac{d_{1}}{2}}C_{1}(M)L_{\frac{1}{2},d_{1}}|\Omega|a^{d_{1}}\lambda^{\frac{d_{1}+1}{2}}+(\frac{d_{1}+2}{d_{1}})^{\frac{d_{1}}{2}}C_{d_{1}}|\Omega|a^{d_{1}}\lambda^{\frac{d_{1}}{2}}.$
and similarly we can control the last term.
[…]
5.4. An abstract extension
As we have seen, although the upper bound given by (5.2) is a bit weaker than (2.6), the extra term 1 can be controlled. Of course one may replace 1 by other number.
More generally, one may start with two increasing sequence
$0<s_{1}\leq s_{2}\leq\cdots+\infty,\qquad t_{1}\leq t_{2}\leq\cdots\to+\infty$
and study the new increasing sequence $\{\nu_{k}(a)\}_{k=1}^{\infty}=\{a^{-2}s_{m}+t_{n}\}$. As usual we will denote
$\mathcal{N}_{(s_{k})}(\lambda)=\#\{k|s_{k}\leq\lambda\}$
and likewise for $\mathcal{N}_{(t_{k})}(\lambda)$. By using the same idea and modifying the proof above slightly, it is easy to prove
Theorem 5.1.
Suppose there exist constants $V_{t},B_{1},B_{2}>0,d\geq 2$ such that
(5.4) $\mathcal{N}_{(t_{k})}(\lambda)\leq V_{t}C_{d}\lambda^{\frac{d}{2}}+B_{1}\lambda^{\frac{d-1}{2}}+B_{2},\qquad\forall\lambda,$
and suppose either $s_{k}=\pi^{2}k^{2}(k\geq 1)$ (in which case we take $V_{s}=1$, $d^{\prime}=1$ below), or there exist $V_{s}>0$ and $d^{\prime}\geq 2$ such that
$\sum_{s_{k}<\lambda}(\lambda-s_{k})<L_{1,{d^{\prime}}}V_{s}\lambda^{\frac{d^{\prime}}{2}+1},\qquad\forall\lambda>0,$
and there exist $C^{\prime}>0$ and $C_{s}>0$ such that for all $\lambda>C_{s}$,
$\sum_{s_{k}<\lambda}(\lambda-s_{k})\leq L_{1,d^{\prime}}V_{s}\lambda^{\frac{d^{\prime}}{2}+1}-C^{\prime}\lambda^{\frac{d^{\prime}+1}{2}},$
then there exists $a_{0}>0$ such that for any $0<a<a_{0}$,
$\nu_{k}(a)\geq\frac{4\pi^{2}}{(\omega_{d+d^{\prime}}a^{d^{\prime}}V_{s}V_{t})^{\frac{2}{d+d^{\prime}}}}k^{\frac{2}{d+d^{\prime}}},\qquad\forall k\geq 1.$
Similarly one may write down an abstract version that extends the results for the Neumann eigenvalues above, in which case one may relax the condition on $\mathcal{N}_{(t_{k})}(\lambda)$ to
(5.5) $\mathcal{N}_{(t_{k})}(\lambda)\geq V_{t}C_{d}\lambda^{\frac{d}{2}}-B_{1}\lambda^{\frac{d-1}{2}},\qquad\forall\lambda>0,$
and pose suitable conditions on $(s_{k})$ (including a Szegö-Weinberger type condition on $s_{1}$).
As a consequence, we could get a bunch of eigenvalue problems that satisfies Pólya inequalities. For example, let $(M,g)$ be a compact Riemannian manifold of dimension $d\geq 2$, with piecewise smooth boundary $\partial M$. Let $(H)$ be certain boundary condition so that the Laplace-Beltrami operator on $(M,g)$ has discrete spectrum. As usual we denote the corresponding eigenvalue counting function by $\mathcal{N}^{(H)}_{M}(\lambda)$. Then we have
Theorem 5.2.
Let $\Omega\subset\mathbb{R}^{d_{1}}$ be a bounded domain with $C^{1}$ boundary and consider the product manifold $a\Omega\times M$.
  1. (1)
    If $\mathcal{N}^{(H)}_{M}(\lambda)$ satisfies (5.4), then there exists $a_{0}>0$ (depends on $\Omega$ and $M$) such that for any $0<a<a_{0}$, the eigenvalues of the Laplace-Beltrami operator on $a\Omega\times M$ with the following mixed boundary condition
[…]
    satisfy Pólya’s conjecture (1.2),
  2. (2)
    If $\mathcal{N}^{(H)}_{M}(\lambda)$ satisfies (5.5), then there exists $a_{0}>0$ (depends on $\Omega$ and $M$) such that for any $0<a<a_{0}$, the eigenvalues of the Laplace-Beltrami operator on $a\Omega\times M$ with the following mixed boundary condition
[…]
    satisfy Pólya’s conjecture (1.3).
For example, one may take the condition (H) to be either Dirichlet boundary condition or Neumann boundary condition or Robin ($\frac{\partial f}{\partial\nu}=\rho f$, with bounded $\rho$) boundary condition, and in all these cases the inequalities (5.4) and (5.5) hold. Thus one get many eigenvalue problems whose eigenvalues satisfy Pólya’s conjecture.
6. Two examples with explicit constants
In this section, we give two examples for which one can calculate the constants involved in the proof, and thus give explicit domains/manifolds for which Pólya’s conjecture holds.
We first construct a planar domain $\Omega$ for which we can calculate the constant $C(\Omega)$ in (2.6) explicitly, and thus find out the number $a_{0}$ in Theorem 1.2 for $\Omega$.
Let $S$ be a square with side length 10 and $T$ be an equilateral triangle with side length 1. The domain $\Omega$ is constructed by placing $T$ at the center of one side of $S$, as shown by the picture below:
Note that the angle $\theta=\frac{2\pi}{3}$, which implies that $\Omega$ cannot tile $\mathbb{R}^{2}$. In what follows we prove
Proposition 6.1.
For any $a\leq\frac{1}{4\pi}$, the Dirichlet eigenvalues of $(0,a)\times\Omega$ satisfies Pólya’s conjecture (1.2).
Proof.
By Faber–Krahn’s inequality ([5], [19]),
$\lambda_{1}(\Omega)\geq\frac{4\pi^{2}}{\omega_{2}|\Omega|}>10^{-1}.$
So $\mathcal{N}^{D}_{\Omega}(\lambda)=0$ for $\lambda\leq 10^{-1}$.
Now suppose $\lambda>10^{-1}$. For the square $S$ we have
$\mathcal{N}^{N}_{S}(\lambda)$ =#$=\#\{(m,n)\in\mathbb{Z}_{\geq 0}^{2}|\ m^{2}+n^{2}<\frac{100\lambda}{\pi^{2}}\}$
$\leq\frac{100\lambda}{4\pi}+\frac{20}{\pi}\sqrt{\lambda}+2$
$<\frac{100\lambda}{4\pi}+20\sqrt{\lambda}.$
For the triangle $T$, let $P=\{(x,2x)\}_{x\in\mathbb{R}}\cup\{(2x,x)\}_{x\in\mathbb{R}}\cup\{(x,-x)\}_{x\in\mathbb{R}}$, then by [28, Proposition 3], one has
$\mathcal{N}^{N}_{T}(\lambda)=$ $\frac{1}{6}\#\{(m,n)\in\mathbb{Z}^{2}|\ (m,n)\notin P,3|(m+n),\frac{16\pi^{2}}{27}(m^{2}+n^{2}-mn)<\lambda\}+$
$\frac{1}{3}\#\{(m,n)\in\mathbb{Z}^{2}|\ (m,n)\in P,\frac{16\pi^{2}}{27}(m^{2}+n^{2}-mn)<\lambda\}+\frac{2}{3}.$
Since
#$\#\{(m,n)\in\mathbb{Z}^{2}|\ (m,n)\notin P,3|(m+n),\frac{16\pi^{2}}{27}(m^{2}+n^{2}-mn)<\lambda\}$
$=$ #$\#\{(m,n)\in\mathbb{Z}^{2}|\ (m,n)\notin P,3|(m+n),(m-\frac{n}{2})^{2}+\frac{3n^{2}}{4}<\frac{27\lambda}{16\pi^{2}}\}$
$\leq$ $\frac{1}{3}\#\{(m,n)\in\mathbb{Z}^{2}|\ m^{2}+\frac{3n^{2}}{4}<\frac{27\lambda}{16\pi^{2}}\}+2(\frac{3\sqrt{3}\sqrt{\lambda}}{4\pi}+\frac{3\sqrt{\lambda}}{2\pi})+4$
$<$ $\frac{3\sqrt{3}\lambda}{8\pi}+60\sqrt{\lambda}$
and
#$\#\{(m,n)\in\mathbb{Z}^{2}|\ (m,n)\in P,\frac{16\pi^{2}}{27}(m^{2}+n^{2}-mn)<\lambda\}$
$\leq$ $3\#\{k\in\mathbb{Z}|\ k^{2}<\frac{9\lambda}{16\pi^{2}}\}<\frac{9\sqrt{\lambda}}{4\pi}+3<30\sqrt{\lambda},$
we get
$\mathcal{N}^{N}_{T}(\lambda)<\frac{\sqrt{3}\lambda}{16\pi}+30\sqrt{\lambda}.$
So we get
$\mathcal{N}^{D}_{\Omega}(\lambda)<\mathcal{N}^{N}_{\Omega}(\lambda)\leq\mathcal{N}^{N}_{S}(\lambda)+\mathcal{N}^{N}_{T}(\lambda)\leq\frac{1}{4\pi}(100+\frac{\sqrt{3}}{4})\lambda+50\sqrt{\lambda}.$
In other words, one may take $C(\Omega)=50$ in (2.6). It follows from the proof in §4.2 that for any $a\leq\frac{1}{4\pi}$, all Dirichlet eigenvalues of $(0,a)\times\Omega$ satisfy Pólya’s Conjecture (1.2). ∎
Finally we turn to the Riemannian manifold setting and consider the standard two-sphere $M=S^{2}$. It is well known that the eigenvalues of $(S^{2},g_{0})$ are $k(k+1)$, with multiplicity $12k+$ 1 for all $k\in\mathbb{Z}_{\geq 0}$. It follows that
$\mathcal{N}_{S^{2}}\big{(}k(k+1)\big{)}=k^{2},\qquad\mathcal{N}_{S^{2}}\big{(}k(k+1)+\varepsilon\big{)}=(k+1)^{2}.$
In other words, one can choose $C(S^{2})$ in (5.1) to be 1 and $C_{1}(S^{2})$ in (5.2) to be 1. Plugging into (5.3) and (4.7), we get
Proposition 6.2.
For any $a\leq\frac{\pi}{24}$, the manifold $(0,a)\times S^{2}$ satisfy Pólya’s conjecture (1.2) and (1.3).
Note that in this example, if we take $a$ large,
  • If we take $a>\sqrt{\frac{2}{3}}\pi$, then the first Dirichlet eigenvalue of $(0,a)\times S^{2}$
    $\lambda_{1}\big{(}(0,a)\times S^{2}\big{)}=\frac{\pi^{2}}{a^{2}}<\frac{4\pi^{2}}{({4\pi\omega_{3}}a)^{\frac{2}{3}}},$
  • If we take $\frac{\pi}{\sqrt{2}}\leq a<\sqrt{\frac{2}{3}}\pi$, then the first nonzero Neumann eigenvalue of $(0,a)\times S^{2}$
    $\mu_{1}\big{(}(0,a)\times S^{2}\big{)}=\frac{\pi^{2}}{a^{2}}>\frac{4\pi^{2}}{({4\pi\omega_{3}}a)^{\frac{2}{3}}},$
so Pólya’s inequalities (1.2) and (1.3) will not hold for $(0,a)\times S^{2}$ when $a$ is large.
Remark 6.3.
We remark that in [13, Example 2.D], P. Freitas and I. Salavessa had already observed (from a very simple tiling argument) that $(0,a)\times S^{2}$ satisfies Pólya’s inequalities for $a$ small enough but fails to satisfy Pólya’s inequalities for $a$ large. Our method is more complicated but has the advantage that we could give explicit estimates of $a$ for Pólya’s inequalities to hold. We would like to thank the authors for pointing out this fact to us.